# AANet: Virtual Screening under Structural Uncertainty via Alignment and Aggregation

### v1

@@ Abstract
Virtual screening (VS) is a critical component of modern drug discovery, yet most existing methods—whether physics-based or deep learning-based—are developed around holo protein structures with known ligand-bound pockets. Consequently, their performance degrades significantly on apo or predicted structures such as those from AlphaFold2, which are more representative of real-world early-stage drug discovery, where pocket information is often missing. In this paper, we introduce an alignment-and-aggregation framework to enable accurate virtual screening under structural uncertainty. Our method comprises two core components: (1) a tri-modal contrastive learning module that aligns representations of the ligand, the holo pocket, and cavities detected from structures, thereby enhancing robustness to pocket localization error; and (2) a cross-attention based adapter for dynamically aggregating candidate binding sites, enabling the model to learn from activity data even without precise pocket annotations. We evaluated our method on a newly curated benchmark of apo structures, where it significantly outperforms state-of-the-art methods in blind apo setting, improving the early enrichment factor (EF1%) from 11.75 to 37.19. Notably, it also maintains strong performance on holo structures. These results demonstrate the promise of our approach in advancing first-in-class drug discovery, particularly in scenarios lacking experimentally resolved protein-ligand complexes.
@@ 1 Introduction
However, adapting existing SBVS methods to predicted structures remains a significant challenge. Prior studies kersten_hic_2023 ; diaz-rovira_are_2023 ; holcomb_evaluation_2023 ; gu_evaluation_2024 have shown that docking performance degrades substantially in predict structures. To address this, flexible refinement zhang2023benchmarking and flexible docking techniques zhu_diffbindfr_2024 ; plainer_diffdock-pocket_2023 ; huang_re-dock_2024 have been proposed to adjust local pocket geometries and better accommodate ligands. However, these methods typically rely on pocket annotations derived from holo structure-an assumption that does not hold in realistic apo or predicted settings. Consequently, these methods largely overlook the upstream challenge of binding site identification under structural uncertainty—a critical bottleneck that severely limits docking performance.
@@ 2 Related work
However, most of these methods assume access to high-quality holo structures and overlook the practical challenges posed by predict apo proteins. The accuracy of conventional docking deteriorates significantly in these settings due to the lack of ligand-induced conformational changes diaz-rovira_are_2023 ; holcomb_evaluation_2023 ; gu_evaluation_2024 , although some approaches attempt to mitigate this through flexible modeling zhang2023benchmarking or by leveraging homologous holo structures kersten_hic_2023 . However, DrugCLIP and other DL-based methods exhibits robustness to structural perturbations; Instead, our analysis suggests that it is highly sensitive to the location and quality of the predefined binding pocket in holo structures.
@@ 3.1 Formulating virtual screening under structural uncertainty
SBVS aims to identify bioactive molecules from a candidate library =$\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, given a protein structure $x_{n}\in\mathbb{R}^{3\times N}$. In conventional settings, the protein is provided in holo form, and the binding pocket $P_{l}$ is defined by the spatial neighborhood surrounding a co-crystallized ligand $x_{m}\in\mathbb{R}^{3\times M}$:
@@ 3.2 Disentangling failure modes in virtual screening without holo structures
We evaluated Glide and the DL-based DrugCLIP gao2023drugclip across these settings. As shown in Figure 1 and Table 1, Glide shows a sharp drop even in the apo-exp (oracle) setting, revealing its sensitivity to conformational changes. In contrast, DrugCLIP maintains performance across holo, apo-exp, and apo-pred (oracle), indicating robustness to moderate structural noise. However, its performance declines significantly in the annotated and blind settings, where pocket localization is uncertain. Even slight misalignments in pocket position lead to notable degradation (see Supplementary Figure).
@@ 3.3 Tri-modal contrastive alignment
In addition, given the observation that the detected pockets can substantially deviate from holo pockets, especially when cramped regions within a pocket cause the detected cavity to split into smaller sub-pockets (see Supplementary), we increased the pocket extraction radius from 6 Å to 10 Å while maintaining the same maximum number of atoms by applying atom down-sampling to mitigate this issue.
Hard negative mining from non-binding cavities. We then complement the positive terms with negative sampling to equip the model with the ability to distinguish true binding pockets from a set of candidate cavities. We introduce negative samples by selecting non-binding detected pockets during training. Specifically, half of the detected cavities with low IoU are treated as non-binding $P_{c}^{-}$ and used as negative pairs under the same pairwise-sigmoid loss (Equation 4) with label $z=-1$ for both $(P_{c}^{-},l)$ and $(P_{c}^{-},P_{l})$.
@@ 4.1 Experimental settings
Training. Our model was initialized with Uni-Mol zhou2023uni and fine-tuned on the PDBBind 2020 general set liu2017pdbbind , with all entities overlapping with DUD-E or LIT-PCBA removed. For dynamic aggregation training, binding and activity data from ChEMBL35 mendez_chembl_2018 were filtered and mapped to predicted apo structures from the AlphaFold Protein Structure Database varadi_alphafolddb_2024 . All UniProt entries matching any PDB entity in DUD-E or LIT-PCBA were excluded to avoid data leakage. The model directly predicts binding scores from protein–ligand pairs and is evaluated in a virtual screening setting without any task-specific fine-tuning.
@@ 4.2 Screening from holo to apo settings
Table 1: Performance on DUD-E and LIT-PCBA. Each method is evaluated on three structural subsets: holo, apo (experimental), and apo (predicted), under both annotated and blind settings. Bold numbers indicate the best performance in each dataset–subset configuration. Row colors indicate method type: Docking & Rescoring, Docking-free baseline, and Proposed method.
@@
holo apo-exp (annot) apo-pred (annot) apo-exp (blind) apo-pred (blind) holo apo-exp (annot) apo-pred (annot) apo-exp (blind) apo-pred (blind)
DUD-E (n = 38)
Glide-SP 0.3124 0.1624 0.2129 18.07 8.66 11.05
RTMScore 0.4311 0.1918 0.2077 26.34 11.47 11.74
EquiScore 0.2479 0.1466 0.1644 14.46 8.44 9.89
TankBind 0.2886 0.2996 0.3008 0.3074 0.2930 17.03 18.13 17.76 18.52 17.36
DrugCLIP 0.5157 0.3493 0.3746 0.1926 0.1974 33.70 21.36 22.70 11.75 12.05
AANet 0.6365 0.5866 0.6003 0.5764 0.6232 40.85 38.03 38.46 37.19 40.85
LIT-PCBA (n = 12)
Glide-SP 0.0565 0.0503 0.0323 5.05 3.06 1.42
RTMScore 0.0445 0.0173 0.0205 3.34 0.67 1.16
EquiScore 0.0556 0.0200 0.0511 4.06 1.27 3.24
TankBind 0.0455 0.0383 0.0484 0.0420 0.0407 3.47 2.36 3.43 1.97 2.48
DrugCLIP 0.0690 0.0554 0.0483 0.0210 0.0155 5.96 4.51 2.85 1.54 0.88
AANet 0.0850 0.0730 0.0805 0.0630 0.0715 7.54 5.58 6.64 3.92 5.40
@@
On the both benchmarks, AANet achieves the highest scores in BEDROC and EF1% across holo and both apo settings, and ranks best in AUROC as well (details in Supplementary Appendix). Notably, its performance remains steady across structural conditions, demonstrating superior robustness to pocket uncertainty and backbone variation. In contrast, both docking-based methods (including rescoring) and the docking-free DrugCLIP show substantial performance degradation. TankBind maintains consistent performance likely due to its use of detected pockets during training rather than ligand-defined ones, as well as its large binding region radius (up to 20 Å), which in some cases covers the entire protein when the target is small.
@@ 4.3 Pocket identification with holo ligand
Table 2: Pocket identification performance on COACH420 dataset (n = 433 ligands/pockets on 288 structures) at various distance cutoff thresholds (Å) from pocket center to any ligand heavy atom (DCA). “Top-1” is the fraction of cases where the highest-scoring pocket lies within the given DCA; “Top-n” is the fraction where any of the top $n$ pockets lies within that DCA. The oracle row is shaded in gray; bold entries denote the best among non-oracle methods.
Method Top-1 ($\mathrm{DCA}\leq x\,\text{\AA}$) Top-n ($\mathrm{DCA}\leq x\,\text{\AA}$)
@@ 4.4.1 Ablation on different modules
Table 3: BEDROC scores on DUD-E (n = 38) under different ablation settings. CA = cavity augmentation, NS = negative sampling, EP = enlarged pocket; PA = pocket adapter, PL = pocket label supervision. For PA column, M = argmax pocket adapter, A = cross-attention adapter.
@@
Since AANet was primarily trained and evaluated using pockets detected by Fpocket, one potential concern is whether its performance stems from overfitting to Fpocket-specific biases rather than learning features intrinsic to the protein structure. To address this, we evaluated AANet using alternative pocket detection methods capra_concavity_2009 . As shown in Table 4, AANet maintains a comparable level of performance on experimental apo structures among alternative pocket detection methods, while that on predicted ones shows a slight decline. To further investigate this discrepancy, we re-detected pockets and re-evaluated performance on predicted apo manually truncated low-confidence regions near the binding site (denoted as apo-pred-t). AANet remained stable across detectors on apo-pred-t, suggesting that the observed gap on full CSMs is more likely attributable to differences in robustness of the detection methods to structural inaccuracies, rather than inherent limitations of the model itself. These results indicate that AANet generalizes well across different pocket sources and does not overly rely on artifacts introduced by a specific detector.
@@ A.3 Pocket external shape and ligand shape
Similarity Actives Full set Top 1% (CPML) EF1% BEDROC $(\alpha=80.5$)
@@ B.1 Pseudo code
protein structure coordinates $x_{n}$, co-crystallized ligand coordinates $x_{m}$
$\{\mathcal{C}^{(s)}\}_{s=1}^{S}\leftarrow\text{detector}(x_{n})$triangleright each cavity: a set of alpha-sphere centers
@@
     $R_{c}^{(s)}\leftarrow\{\,r\in\text{residues}\mid\exists\,x\in r,\ \exists\,c\in\mathcal{C}^{(s)},\ \|x-c\|\leq d\,\}$triangleright residue-level pocket
@@
     $IoU(P_{l},P_{c})=\frac{|P_{l}\cap P_{c}|}{|P_{l}\cup P_{c}|}$triangleright compute overlap ratio between holo pocket and candidate
     $P_{c}\;\leftarrow\;\{P_{c}^{(s)}\}_{s=1}^{S}\;\text{with}\;IoU(P_{l},P_{c})\geq\tau_{\text{pos}}$triangleright sample one positive cavity
     $\{P_{c}^{-}\}\;\leftarrow\;\text{random half of }\bigl{\{}P_{c}^{(s)}:IoU(P_{l},P_{c}^{(s)})\leq\tau_{\text{neg}}\bigr{\}}$triangleright sample hard negatives
@@
         $\ell^{(s)}\leftarrow\langle e_{l},\,e_{c}^{(s)}\rangle\;/\;t$triangleright compute dot-product logits
@@
     $a^{(s)}\leftarrow\exp(\ell^{(s)})\big{/}\sum_{j=1}^{S}\exp(\ell^{(j)})\quad\forall s$triangleright softmax attention
     $\tilde{e}_{c}\leftarrow\sum_{s=1}^{S}a^{(s)}\,e_{c}^{(s)}$triangleright aggregate cavities
@@
         $s_{s}=1$ for the true cavity, otherwise 00
@@
     $\mathcal{L}_{\mathrm{KL}}^{\prime}\leftarrow\sum_{s=1}^{S}s_{s}\,\log\!\bigl{(}\tfrac{s_{s}}{a^{(s)}}\bigr{)}$triangleright attention supervision
@@
B.2 Benchmark details
@@
ADRB2 P07550 9chv A, B, C Non-redundant protein complex retained
@@
IDH1 O75874 1t0l A, B Ligand in holo (6ADG) binds at the interface of two homologous chains
@@
MAPK1 P28482 4iz7 A, B Non-redundant protein complex retained
@@
TP53 P04637 1kzy A, C Non-redundant protein complex retained
@@ B.3 Hyperparameters
Loss logit bias $-10$
@@
B.4 Experiments compute resources
@@
B.5 Metrics
We evaluate virtual screening performance using three complementary measures that capture both overall ranking quality and early-retrieval effectiveness.

### v2

@@ Abstract
Virtual screening (VS) is a critical component of modern drug discovery, yet most existing methods—whether physics-based or deep learning-based—are developed around holo protein structures with known ligand-bound pockets. Consequently, their performance degrades significantly on apo or predicted structures such as those from AlphaFold2, which are more representative of real-world early-stage drug discovery, where pocket information is often missing. In this paper, we introduce an alignment-and-aggregation framework to enable accurate virtual screening under structural uncertainty. Our method comprises two core components: (1) a tri-modal contrastive learning module that aligns representations of the ligand, the holo pocket, and cavities detected from structures, thereby enhancing robustness to pocket localization error; and (2) a cross-attention based adapter for dynamically aggregating candidate binding sites, enabling the model to learn from activity data even without precise pocket annotations. We evaluated our method on a newly curated benchmark of apo structures, where it significantly outperforms state-of-the-art methods in blind apo setting, improving the early enrichment factor (EF1%) from 11.75 to 37.19. Notably, it also maintains strong performance on holo structures. These results demonstrate the promise of our approach in advancing first-in-class drug discovery, particularly in scenarios lacking experimentally resolved protein-ligand complexes. Our implementation is publicly available at https://github.com/Wiley-Z/AANet.
@@ 1 Introduction
However, adapting existing SBVS methods to predicted structures remains a significant challenge. Prior studies kersten_hic_2023 ; diaz-rovira_are_2023 ; holcomb_evaluation_2023 ; gu_evaluation_2024 have shown that docking performance degrades substantially in predict structures. To address this, flexible refinement zhang2023benchmarking and flexible docking techniques zhu_diffbindfr_2024 ; plainer_diffdock-pocket_2023 ; huang_re-dock_2024 ; morehead2025flowdock have been proposed to adjust local pocket geometries and better accommodate ligands. However, these methods typically rely on pocket annotations derived from holo structure-an assumption that does not hold in realistic apo or predicted settings. Consequently, these methods largely overlook the upstream challenge of binding site identification under structural uncertainty—a critical bottleneck that severely limits docking performance.
@@ 2 Related work
Nonetheless, most of these methods assume access to high-quality holo structures and overlook the practical challenges posed by predict apo proteins. The accuracy of conventional docking deteriorates significantly in these settings due to the lack of ligand-induced conformational changes diaz-rovira_are_2023 ; holcomb_evaluation_2023 ; gu_evaluation_2024 , although some approaches attempt to mitigate this through flexible modeling zhang2023benchmarking or by leveraging homologous holo structures kersten_hic_2023 . However, DrugCLIP and other DL-based methods exhibits robustness to structural perturbations; Instead, our analysis suggests that it is highly sensitive to the location and quality of the predefined binding pocket in holo structures.
@@ 3.1 Formulating virtual screening under structural uncertainty
SBVS aims to identify bioactive molecules from a candidate library $\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, given a protein structure $x_{n}\in\mathbb{R}^{3\times N}$. In conventional settings, the protein is provided in holo form, and the binding pocket $P_{l}$ is defined by the spatial neighborhood surrounding a co-crystallized ligand $x_{m}\in\mathbb{R}^{3\times M}$:
@@ 3.2 Disentangling failure modes in virtual screening without holo structures
We evaluated Glide and the DL-based DrugCLIP gao2023drugclip across these settings. As shown in Figure 1 and Table 1, Glide shows a sharp drop even in the apo-exp (oracle) setting, revealing its sensitivity to conformational changes. In contrast, DrugCLIP maintains performance across holo, apo-exp, and apo-pred (oracle), indicating robustness to moderate structural noise. However, its performance declines significantly in the annotated and blind settings, where pocket localization is uncertain. Even slight misalignments in pocket position lead to notable degradation (see Appendix A.1).
@@ 3.3 Tri-modal contrastive alignment
In addition, given the observation that the detected pockets can substantially deviate from holo pockets, especially when cramped regions within a pocket cause the detected cavity to split into smaller sub-pockets (see Appendix A.2), we increased the pocket extraction radius from 6 Å to 10 Å while maintaining the same maximum number of atoms by applying atom down-sampling to mitigate this issue.
Hard negative mining from non-binding cavities. We then complement the positive terms with negative sampling to equip the model with the ability to distinguish true binding pockets from a set of candidate cavities. We introduce negative samples by selecting non-binding detected pockets during training. Specifically, half of the detected cavities with low IoU are treated as non‐binding $P_{c}^{-}$ and used as negative pairs under the same pairwise‐sigmoid loss (Equation 4) with label z=1z=-1 for both $(P_{c}^{-},l)$ and $(P_{c}^{-},P_{l})$.
@@ 4.1 Experimental settings
Training. Our model was initialized with Uni-Mol zhou2023uni and fine-tuned on the PDBBind 2020 general set liu2017pdbbind , with all entities overlapping with DUD-E or LIT-PCBA removed. For dynamic aggregation training, binding and activity data from ChEMBL35 mendez_chembl_2018 (pre-processing described in B.2) were filtered and mapped to predicted apo structures from the AlphaFold Protein Structure Database varadi_alphafolddb_2024 . All UniProt entries matching any PDB entity in DUD-E or LIT-PCBA were excluded to avoid data leakage. The model directly predicts binding scores from protein–ligand pairs and is evaluated in a virtual screening setting without any task-specific fine-tuning.
@@ 4.2 Screening from holo to apo settings
Table 1: Performance on DUD‑E and LIT‑PCBA. Each method is evaluated on three structural subsets: holo, apo (experimental), and apo (predicted), under both annotated and blind settings. Bold numbers indicate the best performance in each dataset–subset configuration. Row colors indicate method type: Docking & Rescoring, Docking-free baseline, and Proposed method.
@@
holo
apo-exp
apo-pred
apo-exp
apo-pred
holo
apo-exp
apo-pred
apo-exp
apo-pred
DUD‑E (n = 38)
[…]
LIT‑PCBA (n = 12)
[…]
@@
On the both benchmarks, AANet achieves the highest scores in BEDROC and EF1% across holo and both apo settings, and ranks best in AUROC as well (details in Appendix C.2). Notably, its performance remains steady across structural conditions, demonstrating superior robustness to pocket uncertainty and backbone variation. In contrast, both docking-based methods (including rescoring) and the docking-free DrugCLIP show substantial performance degradation. TankBind maintains consistent performance likely due to its use of detected pockets during training rather than ligand-defined ones, as well as its large binding region radius (up to 20 Å), which in some cases covers the entire protein when the target is small.
@@ 4.3 Pocket identification with holo ligand
Table 2: Pocket identification performance on COACH420 dataset (n = 433 ligands/pockets on 288 structures) at various distance cutoff thresholds (Å) from pocket center to any ligand heavy atom (DCA). “Top‑1” is the fraction of cases where the highest‑scoring pocket lies within the given DCA; “Top‑n” is the fraction where any of the top $n$ pockets lies within that DCA. The oracle row is shaded in gray; bold entries denote the best among non‑oracle methods.
Method Top‑1 ($\mathrm{DCA}\leq x\,\text{\AA }$) Top‑n ($\mathrm{DCA}\leq x\,\text{\AA }$)
@@ 4.4.1 Ablation on different modules
Table 3: BEDROC scores on DUD‑E (n = 38) under different ablation settings. CA = cavity augmentation, NS = negative sampling, EP = enlarged pocket; PA = pocket adapter, PL = pocket label supervision. For PA column, M = argmax pocket adapter, A = cross-attention adapter.
@@
Since AANet was primarily trained and evaluated using pockets detected by Fpocket, one potential concern is whether its performance stems from overfitting to Fpocket-specific biases rather than learning features intrinsic to the protein structure. To address this, we evaluated AANet using alternative pocket detection methods capra_concavity_2009 . As shown in Table 4, AANet maintains a comparable level of performance on experimental apo structures among alternative pocket detection methods, while that on predicted ones shows a slight decline. To further investigate this discrepancy, we re-detected pockets and re-evaluated performance on predicted apo manually truncated low-confidence regions near the binding site (denoted as apo-pred-t). AANet remained stable across detectors on apo-pred-t, suggesting that the observed gap on full predicted apo structures is more likely attributable to differences in robustness of the detection methods to structural inaccuracies, rather than inherent limitations of the model itself. These results indicate that AANet generalizes well across different pocket sources and does not overly rely on artifacts introduced by a specific detector.
@@ 5 Conclusion
Acknowledgements
This work is supported by Beijing Academy of Artificial Intelligence and Beijing Frontier Research Center for Biological Structure Fundings.
@@ A.3 Pocket external shape and ligand shape
Similarity Actives Full set Top 1% (CPML) EF1%
BEDROC
$(\alpha=80.5$)
@@ B.1 Pseudo code
protein structure coordinates $x_{n}$, co‐crystallized ligand coordinates $x_{m}$
$\{\mathcal{C}^{(s)}\}_{s=1}^{S}\leftarrow\text{detector}(x_{n})$ $\triangleright$ each cavity: a set of alpha-sphere centers
@@
  $R_{c}^{(s)}\leftarrow\{\,r\in\text{residues}\mid\exists\,x\in r,\ \exists\,c\in\mathcal{C}^{(s)},\ \|x-c\|\leq d\,\}$ $\triangleright$ residue-level pocket
@@
  $IoU(P_{l},P_{c})=\frac{|P_{l}\cap P_{c}|}{|P_{l}\cup P_{c}|}$ $\triangleright$ compute overlap ratio between holo pocket and candidate
  $P_{c}\;\leftarrow\;\{P_{c}^{(s)}\}_{s=1}^{S}\;\text{with}\;IoU(P_{l},P_{c})\geq\tau_{\text{pos}}$ $\triangleright$ sample one positive cavity
  $\{P_{c}^{-}\}\;\leftarrow\;\text{random half of }\bigl\{P_{c}^{(s)}:IoU(P_{l},P_{c}^{(s)})\leq\tau_{\text{neg}}\bigr\}$ $\triangleright$ sample hard negatives
@@
   $\ell^{(s)}\leftarrow\langle e_{l},\,e_{c}^{(s)}\rangle\;/\;t$ $\triangleright$ compute dot-product logits
@@
  $a^{(s)}\leftarrow\exp(\ell^{(s)})\big/\sum_{j=1}^{S}\exp(\ell^{(j)})\quad\forall s$ $\triangleright$ softmax attention
  $\tilde{e}_{c}\leftarrow\sum_{s=1}^{S}a^{(s)}\,e_{c}^{(s)}$ $\triangleright$ aggregate cavities
@@
   $s_{s}=1$ for the true cavity, otherwise 0
@@
  $\mathcal{L}_{\mathrm{KL}}^{\prime}\leftarrow\sum_{s=1}^{S}s_{s}\,\log\!\bigl(\tfrac{s_{s}}{a^{(s)}}\bigr)$ $\triangleright$ attention supervision
@@
B.2 ChEMBL data processing
Molecule filtering:
  • Removed salts or kept only the largest fragment
  • Kept molecules with molecular weight between [100, 800]
  • Removed those containing atoms other than [H, C, N, O, F, Cl, Br, I, S, P, B, Se]
  • Filtered out molecules with unbranched long chains containing 6 atoms or more.
Activity filtering: We retained activity records with:
  • Confidence score = 9
  • Assay type = functional or binding
  • Standard type $\in$ [’Ki’, ’IC50’, ’Kd’, ’EC50’, ’ED50’, ’AC50’, ’XC50’]
  • Values converted to molar units and filtered to lie within the –log10 range [5, 12]
Target mapping: Protein targets were mapped to UniProt IDs and matched with AlphaFold structures. Targets without available predicted structures were discarded.
Contrastive supervision masking: Each known protein–ligand activity pair was recorded using its UniProt ID and InChI-key. During training, we masked non-diagonal entries that correspond to active (positive) pairs to ensure that true positive pairs are never treated as negatives, even if they are not aligned in the current batch. This masking strategy ensures correct supervision and avoids misleading the model during contrastive learning.
B.3 Benchmark details
@@
ADRB2 P07550 9chv A, B, C
Non-redundant protein
complex retained
@@
IDH1 O75874 1t0l A, B
Ligand in holo (6ADG) binds
at the interface of two
homologous chains
@@
MAPK1 P28482 4iz7 A, B
Non-redundant protein
complex retained
@@
TP53 P04637 1kzy A, C
Non-redundant protein
complex retained
@@ B.4 Hyperparameters
Loss logit bias 10-10
@@
B.5 Experiments compute resources
@@
B.6 Metrics
We evaluate virtual screening performance using three complementary measures that capture both overall ranking quality and early‐retrieval effectiveness.